OpenStudy (complex):

(√−4) / (√−3)(√−7) I keep coming up with 3.055 for the imaginary number (its not right). I know the real number is 0.

OpenStudy (anonymous):

figure it out then

OpenStudy (complex):

trying..

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

what grade are you in

OpenStudy (campbell_st):

so is the question 1.\[\frac{\sqrt{-4}}{\sqrt{-3}\times \sqrt{-7}}\] or 2. \[\frac{\sqrt{-4}}{\sqrt{-3}} \times \sqrt{-7}\]

OpenStudy (anonymous):

$$\frac{\sqrt{−4}}{\sqrt{−3}\sqrt{−7}}=\frac2{\sqrt{21}}i$$

OpenStudy (anonymous):

$$3.055i\approx2\sqrt{7/3}\ i=\frac{\sqrt{-4}}{\sqrt{-3}}\sqrt{-7}$$